### v2

# Bi-Hamiltonian structures of WDVV-type

Abstract
We study a class of nonlinear PDEs that admit the same bi-Hamiltonian structure as WDVV equations: a Ferapontov-type first-order Hamiltonian operator and a homogeneous third-order Hamiltonian operator in a canonical Doyle–Potëmin form, which are compatible. Using various equivalence groups, we classify such equations in two-component and three-component cases. In a four-component case, we add further evidence to the conjecture that there exists only one integrable system of the above type. Finally, we give an example of the six-component system with required bi-Hamiltonian structure. To streamline the symbolic computation we develop an algorithm to find the aforementioned Hamiltonian operators, which includes putting forward a conjecture on the structure of the metric parameterising the first-order Hamiltonian operator.
∗∗Dipartimento di Matematica e Fisica “E. De Giorgi”,
Università del Salento
and Sezione INFN di Lecce
via per Arnesano, 73100 Lecce, Italy
 Institute of Mathematics of NAS of Ukraïne,
3 Tereshchenkivska Str., 01024 Kyiv, Ukraïne
e-mails: stanislav.opanasenko@le.infn.it
raffaele.vitolo@unisalento.it
1 Introduction
1.1 Integrable Systems and WDVV equations
In the area of infinite-dimensional integrable systems, bi-Hamiltonian systems of partial differential equations (PDEs) play a central role [7, 29, 30, 34, 35, 47]. The presence of two independent compatible Hamiltonian structures ensures, under certain mild hypotheses [43], the existence of infinite sequences of commuting conserved quantities and commuting symmetries, thus mimicking the Liouville integrability of the finite-dimensional case.
It was proved by B. Dubrovin that the solutions of Witten–Dijkgraaf–Verlinde–Verlinde (WDVV) equations can be put in correspondence with a large class of bi-Hamiltonian structures (see [14]), thus providing a bridge between Topological Field Theories and Integrable Systems. The consequences in mathematics have been far-reaching, as solutions of WDVV equations yield Dubrovin–Frobenius manifolds, which nowadays is an active research topic.
In dimension $N$, $F=F(t^{1},\dots,t^{N})$, WDVV equations is a nonlinear overdetermined system of PDEs,
$\eta^{\lambda\mu}F_{\lambda\alpha\beta}F_{\mu\nu\gamma}=\eta^{\lambda\mu}F_{\lambda\alpha\nu}F_{\mu\beta\gamma},\quad\text{where}\quad F_{\alpha\beta\gamma}:=\frac{\partial^{3}F}{\partial t^{\alpha}\partial t^{\beta}\partial t^{\gamma}}.$
The inverse $(\eta_{\alpha\beta})$ of the constant symmetric nondegenerate matrix $(\eta^{\alpha\beta})$ is by definition $\eta_{\alpha\beta}:=F_{1\alpha\beta}$. The WDVV equations are equivalent to the requirement of associativity of a product operation with structure constants $c^{\alpha}_{\beta\gamma}=\eta^{\alpha\nu}F_{\nu\beta\gamma}$, and therefore they are sometimes also called associativity equations. Note that the requirements on $F$ completely specify its dependence on $t^{1}$ up to second degree polynomials,
$F=\frac{1}{6}\eta_{11}(t^{1})^{3}+\frac{1}{2}\sum_{k>1}\eta_{1k}t^{k}(t^{1})^{2}+\frac{1}{2}\sum_{k,s>1}\eta_{sk}t^{s}t^{k}t^{1}+f(t^{2},\ldots,t^{N}).$
The fact that systems whose solutions yield integrable systems can themselves be formulated as integrable systems is a common phenomenon in the field of integrable systems. Thus, it was proved [17] that a system of WDVV equations, after being reformulated as a quasilinear first-order system of PDEs in conservative form,
$u^{i}_{t}=(V^{i})_{x},$ (1)
admits a bi-Hamiltonian formulation:
[…]
(we use Einstein’s summation convention throughout the paper). Here $u=(u^{1},\dots,u^{n})$ is the vector of field variables, $(t,x)$ are the independent variables and $V^{i}$’s are smooth functions of the dependent variables. Apart from a conservative form, the quasilinear system we consider have another two important properties: non-diagonalisability and linear degeneracy. Recall that a quasilinear system $u^{i}_{t}=V^{i}_{j}u^{j}_{x}$ is called diagonalisable if there exists a change of coordinates $r^{i}=r^{i}(u)$ such that the transformed system is diagonal, $r^{i}_{t}=\tilde{V}^{i}r^{i}_{x}$, and linearly degenerate if $L_{p}v=0$ for every pair $(v,p)$, where $p$ is the right eigenvector of the matrix $(V^{i}_{j})$ corresponding to its eigenvalue $v$ ($L$ is the Lie derivative). The pair of compatible Hamiltonian operators $A_{1}$, $A_{2}$ (compatibility means $A_{1}+\lambda A_{2}$ is a Hamiltonian operator for any $\lambda\in\mathbb{R}$ as well) was different from the pair of compatible Hamiltonian operators which is determined by a solution of WDVV equations. Bi-Hamiltonianity was further confirmed for other simple cases of WDVV equations in [25, 26]. After some years, advances both in the theory [23] and in Symbolic Computations [45, 44, 46, 27] led to the discovery of new bi-Hamiltonian structures for more complicated WDVV equations [39, 40, 44, 45].
1.2 Bi-Hamiltonian structures for WDVV systems
A matrix differential operator $A^{ij}=a^{ij\sigma}(u_{\tau})D_{\sigma}$, where $D_{\sigma}=D_{x}^{\sigma}$, $u_{\tau}$ stands for a finite collection of derivatives of the dependent variables and $\sigma$, $\tau\in\mathbb{N}_{0}$, is called Hamiltonian if the corresponding bracket
[…]
is Poisson. More precisely, skew-symmetry of the bracket is equivalent to skew-adjointness of $A$, and the Jacobi property is equivalent to differential conditions on the coefficients of $A$ that have an intrinsic formulation, the vanishing of the Schouten bracket: $[A,A]=0$. In turn, compatibility of two Hamiltonian operators $A_{1}$, $A_{2}$ is equivalent to the vanishing of the Schouten bracket, $[A_{1},A_{2}]=0$.
Both Hamiltonian operators $A_{1}$, $A_{2}$ admitted by a WDVV system are homogeneous, where homogeneity is defined with respect to the grading $\deg{D_{x}}=1$. Such operators were introduced in [13, 15] as a family of operators that is invariant with respect of diffeomorphisms of the space of field variables.
In particular, the operator $A_{1}$ a first-order homogeneous Hamiltonian operator $A_{1}=P$ of Ferapontov type,
$P^{ij}=g^{ij}\mathrm{D}_{x}+\Gamma^{ij}_{k}u^{k}_{x}+c^{\alpha\beta}w^{i}_{\alpha h}u^{h}_{x}\mathrm{D}_{x}^{-1}w^{j}_{\beta k}u^{k}_{x},$ (3)
and an operator $A_{2}$ is the third-order Hamiltonian operator $A_{2}=R$,
$R^{ij}=\mathrm{D}_{x}(f^{ij}\mathrm{D}_{x}+c^{ij}_{k}u^{k}_{x})\mathrm{D}_{x},$ (4)
which is compatible with $P$. All coefficient functions $g^{ij}$, $f^{ij}$, $\Gamma^{ij}_{k}$, $c^{ij}_{k}$ and $w^{i}_{\alpha h}$ above are functions of the field variables $u$ only, and $c^{\alpha\beta}$ are constants.
For example, in the simplest case, $N=3$ and
$\eta=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}$
the WDVV equations reduce to the form
$f_{ttt}=f_{txx}^{2}-f_{xxx}f_{ttx}.$ (5)
Following [19], we introduce new dependent variables $u^{1}=f_{xxx}$, ${u^{2}=f_{txx}}$, $u^{3}=f_{ttx}$. Then, the equation (5) can be rewritten in the form (1) as
$u^{1}_{t}=u^{2}_{x},\quad u^{2}_{t}=u^{3}_{x},\quad u^{3}_{t}=((u^{2})^{2}-u^{1}u^{3})_{x}.$ (6)
It was found in [17] that the system (6) is bi-Hamiltonian with the local (i.e. $c^{\alpha\beta}=0$) first-order homogeneous operator $P$ and the third-order homogeneous operator $R$,
$P=\begin{pmatrix}-\frac{3}{2}\mathrm{D}_{x}&\frac{1}{2}\mathrm{D}_{x}u^{1}&\mathrm{D}_{x}u^{2}\\ \frac{1}{2}u^{1}\mathrm{D}_{x}&\frac{1}{2}(\mathrm{D}_{x}u^{2}+u^{2}\mathrm{D}_{x})&\frac{3}{2}u^{3}\mathrm{D}_{x}+u^{3}_{x}\\ u^{2}\mathrm{D}_{x}&\frac{3}{2}\mathrm{D}_{x}u^{3}-u^{3}_{x}&((u^{2})^{2}-u^{1}u^{3})\mathrm{D}_{x}+\mathrm{D}_{x}((u^{2})^{2}-u^{1}u^{3})\end{pmatrix},$
$R=\mathrm{D}_{x}\begin{pmatrix}0&0&\mathrm{D}_{x}\\ 0&\mathrm{D}_{x}&-\mathrm{D}_{x}u^{1}\\ \mathrm{D}_{x}&-u^{1}\mathrm{D}_{x}&\mathrm{D}_{x}u^{2}+u^{2}\mathrm{D}_{x}+u^{1}\mathrm{D}_{x}u^{1}\end{pmatrix}\mathrm{D}_{x},$
The corresponding Hamiltonian densities are
$h_{P}=u^{3}\quad\text{and}\quad h_{R}=-\frac{1}{2}u^{1}\left(\mathrm{D}_{x}^{-1}u^{2}\right)^{2}-(\mathrm{D}_{x}^{-1}u^{2})(\mathrm{D}_{x}^{-1}u^{3}).$
Remark 1.
We stress that the type of the bi-Hamiltonian pair for WDVV equations is different from the bi-Hamiltonian pair determined by a solution of WDVV equations. Indeed, the latter is constituted by two compatible local homogeneous Hamiltonian operators of order one [12, 14].
1.3 Problem and results
Definition 2.
We define a bi-Hamiltonian structure of WDVV-type to be a pair of compatible Hamiltonian operators $P$, $R$ as in (3), (4), respectively.
We say a quasilinear first-order system of PDEs in conservative form (1) to be a bi-Hamiltonian system of WDVV-type if it is endowed with a bi-Hamiltonian structure of WDVV-type.
The problem: In this paper, we aim at classifying (when possible) bi-Hamiltonian equations of WDVV-type, and at the same time at introducing new bi-Hamiltonian equations of WDVV-type, which are not necessarily related with WDVV equations.
As this task presents several theoretical and computational challenges, our paper contains both the theoretical advances and the computational algorithms that made this research possible.
The main results of the paper are listed below, following the number of unknown functions $u^{i}$. Note that we discard linear bi-Hamiltonian systems of WDVV type as we regard them as trivial.
$n=2$
We have an affine classification of bi-Hamiltonian structures of WDVV-type, according to which there exist two classes of nonlinear bi-Hamiltonian systems of WDVV type. Both cases turn out to be linearisable if we enlarge the group action to the group of projective reciprocal transformations that preserve $t$.
$n=3$
In this case, under the action of the group of projective reciprocal transformations that preserve $t$, we have five nontrivial cases. Three of them are particular types of WDVV equations known in the literature, while two of them are new. In order to compute the leading coefficient of the operator $P$ we put forward Conjecture 4 in Subsection 2.3 that proved to be true in all known cases.
Then, under the action of the full group of projective reciprocal transformations, it can be shown that all nontrivial cases reduce to one, the simplest WDVV equation. Unfortunately, the corresponding transformation is practically impossible to find with current computational tools. This implies that our methods, which allow to find the bi-Hamiltonian structure, are still highly valuable.
$n=4$
In this case, systems with third-order Hamiltonian structures have been classified, however, only one of them is known to be integrable. The system is obtained within the geometric theory of linearly degenerate systems in the Temple class [5] (no relationship with WDVV). We found a new bi-Hamiltonian structure of WDVV-type for this system. In some other interesting cases we proved that such structures do not exist, although we cannot guarantee it in general.
$n=6$
We proved that two commuting systems of the type (1) are bi-Hamiltonian systems of WDVV-type. These are two integrable systems that arise from integrability conditions of a certain class of Lagrangians [24]. The systems are not related with WDVV equations; this, together with the previous item, proves that the class of bi-Hamiltonian systems of WDVV-type does not reduce to WDVV equations only.
It is worth to compare bi-Hamiltonian structures in this paper with another family of bi-Hamiltonian structures that has been introduced in [28]. Such structures are of the form $A_{1}=P$, $A_{2}=Q+\epsilon^{2}R$, and are said to be bi-Hamiltonian structures of KdV-type. The operators $(P,Q,R)$ are a compatible triple of homogeneous Hamiltonian operators, with $P$, $Q$ being of the form (3) and $R$ being a higher order homogeneous Hamiltonian operator. Note that $\epsilon$ is introduced as a perturbative parameter. For example, the triple
$P=D_{x},\quad Q=\frac{2}{3}u\mathrm{D}_{x}+\frac{1}{3}u_{x},\quad R=\mathrm{D}_{x}^{3}$
provides a bi-Hamiltonian structure for the KdV equation. Many integrable systems have this form (see [28] for a detailed list of examples and references).
Remark 3.
When comparing bi-Hamiltonian structures of KdV and WDVV type, it is evident that we can treat the bi-Hamiltonian structures of WDVV type as a singular case of bi-Hamiltonian structures of KdV type. The ‘singularity’ of the bi-Hamiltonian pair goes in opposite direction with respect to the usual dispersionless limit $\epsilon\to 0$.
The results that we obtained in this paper show that the class of bi-Hamiltonian structures of WDVV type is rich and interesting, and it deserves further investigation.
1.4 Computational problems and their solutions
Generally speaking, finding Hamiltonian operators for systems of PDEs of the type (1) in high ($n\geq 3$) dimension and proving their compatibility is not an easy task. The computational problems to be solved are summarised below.
  • Finding a Hamiltonian formulation (2) through a first-order operator $P$ as in (3) amounts to solving a complicated system of nonlinear PDEs. Such a system is not easily solvable even when $n=3$.
  • On the other hand, third-order operators are classified [21, 22] in low dimensions, as well as the systems of the form (1) that admit a corresponding Hamiltonian formulation [23]. Hence we can start from a third-order operator $R$ from the classification and the corresponding systems, and find first-order Hamiltonian formulations (2).
  • First-order operators $P$ for the above systems of PDEs are found by solving the system mentioned above, with the additional requirement of compatibility: $[P,R]=0$.
Compatibility is the major computational problem here. First of all, until recently, there did not exist a way to bring Schouten bracket between nonlocal operators in a canonical form, and require its vanishing. In [10] an algorithm was presented. However, finding a minimal set of conditions which are equivalent to $[P,R]=0$ is the problem that is still unresolved.
In [9] the above algorithm was implemented in the computer algebra systems Maple, Mathematica and Reduce. Thus, in principle, we can compute $[P,R]$ and require its vanishing if one of the two operators is unknown. However, the computational complexity is sometimes overwhelming, even for dedicated compute servers, and the main problem is to keep the complexity within the limit of what our computers can do.
The requirement $[P,R]=0$ for unknown $P$ leads to a complicated overdetermined nonlinear system of PDEs that we are able to solve only in dimension $n=2$ and in the simplest case in dimension $3$. Working in Maple computer algebra system, we used rifsimp to reduce the equations to an involutive system with a known dimension of finite-dimensional solution space and then pdsolve to solve it.
For more general calculations, we used Conjecture 4; it yields a condition that we always verified in concrete compatible operators $P$ and $R$. Such a condition is very likely to be an important part of the compatibility conditions, and brings again the problem of finding $P$ to a system of algebraic equations, which becomes manageable also in high dimensions.
However, that is not all: we still need to check that $[P,R]=0$, although at this point that is only a straightforward computation as $P$ and $R$ are completely specified. Not all such calculations can be performed on a modern laptop: primarily, huge amounts of RAM are needed as the formula of the Schouten bracket (see Subsection 2.3) and the algorithm for bringing the expressions to a canonical form [10] imply the calculation of iterated derivatives of large rational expressions, leading to expression swell that only at the very end simplifies to zero (if that is the case).
In particular, the largest successful direct calculations have been performed for the calculations of $[P,Q]$ in the case $R^{(2)}$ ($n=3$) ($15$GB of RAM and $90$ hours of computing time) and in the case $n=4$. In the case $R^{(1)}$ ($n=3$) the calculation of $[P,Q]$ failed for lack of RAM. This led to a re-analysis of the algorithm in [10] and a work-around has been found. Since in the computation all denominators are of the form $\sigma^{n}$, where $\sigma$ is the singular variety of the corresponding Monge metric, while $n$ varies, an expression $\sigma^{-1}$ was given a new notation, effectively making all rational expressions polynomials. This modification led to a much reduced RAM consumption in case $R^{(2)}$ (only $2$GB!), and manyfold reduction of computing time, and allowed us to prove compatibility of the Hamiltonian operators $R^{(1)}$ and $P^{(1)}$ using $45$GB of RAM and $54$ hours of computing time.
Surprisingly, the case $n=6$ turned out to be manageable on a laptop with $16$GB of RAM, mostly due to its locality.
All large calculations have been performed on a compute server of the INFN, Section of Lecce. The server has a processor AMD EPYC 7282 and 256GB of RAM. The most important computations that have been programmed by us are available at [37]; we will be happy to help other researchers who are interested in our software or similar computational tasks.
2 A pair of Hamiltonian structures
Our goal is to classify quasilinear first-order systems of evolutionary PDEs of the type (1) that are endowed with a pair of Hamiltonian operators (3) and (4).
2.1 First-order Hamiltonian operator
of Ferapontov type
A Hamiltonian operator $P$,
$P^{ij}=g^{ij}\mathrm{D}_{x}+\Gamma^{ij}_{s}u^{s}_{x}+c^{\alpha\beta}w^{i}_{\alpha s}u^{s}_{x}\mathrm{D}_{x}^{-1}w^{j}_{\beta t}u^{t}_{x},$
is a nonlocal generalisation of a classical Dubrovin–Novikov Hamiltonian operator ($c^{\alpha\beta}=0$), which was introduced by Ferapontov in [16]. The operator $P$ is Hamiltonian if and only if the following properties are fulfilled (we assume the non-degeneracy condition $\det(g^{ij})\neq 0$):
$g^{ij}=g^{ji},\quad g^{ij}_{,k}=\Gamma^{ij}_{k}+\Gamma^{ji}_{k},\quad g^{is}\Gamma^{jk}_{s}=g^{js}\Gamma^{ik}_{s},$ (7a)
$g^{is}w^{j}_{\alpha s}=g^{js}w^{i}_{\alpha s},$ (7b)
$\nabla_{k}w^{i}_{\alpha j}=\nabla_{j}w^{i}_{\alpha k},$ (7c)
$[w_{\alpha},w_{\beta}]=0,$ (7d)
$c^{\alpha\beta}=c^{\beta\alpha},$ (7e)
$R^{ij}_{kl}=c^{\alpha\beta}\Big{(}w^{i}_{\alpha l}w^{j}_{\beta k}-w^{i}_{\alpha k}w^{j}_{\beta l}\Big{)}.$ (7f)
[…]
The number $m$ of the commuting vector fields (and hence the dimension $m$ of the symmetric matrix $(c^{\alpha\beta})$ is not known a priori. However, there is an important computational feature 111 It is not a theorem, but an experimental observation communicated to us by E.V. Ferapontov. of linearly degenerate non-diagonalisable systems is that in low dimension ($n\leq 5$) they only admit two commuting flows of first order, namely the $x$- and $t$-translations, represented by
[…]
All but one of our examples fall within this category of equations, and the one that does not is local, see Section 6, and therefore we shall restrict ourselves to this case hereafter.
The above operator $P$ is clearly invariant with respect to local diffeomorphisms of the dependent variables: $\tilde{u}=\tilde{U}(u)$. Under such transformations, $g^{ij}$ transform as a contravariant $2$-tensor, and $\Gamma^{ij}_{k}$ transform as the contravariant Christoffel symbols of a linear connection. Thus, geometrically the conditions (7) mean that if $(g_{ij})=(g^{ij})^{-1}$, then $\Gamma^{i}_{jk}=-g_{jh}\Gamma^{hi}_{k}$ are the Christoffel symbols of the Levi-Civita connection of the metric $g_{ij}$. Its curvature, expressed via the Kobayashi–Nomizu convention:
$R^{i}_{jkl}=\Gamma^{i}_{lj,k}-\Gamma^{i}_{kj,l}+\Gamma^{i}_{kr}\Gamma^{r}_{lj}-\Gamma^{i}_{lr}\Gamma^{r}_{kj},$
admits the expansion (7f), where
$R^{ijk}_{l}=g^{is}g^{jt}R^{k}_{tsl}=g^{is}(\partial_{l}\Gamma^{jk}_{s}-\partial_{s}\Gamma^{jk}_{l})+\Gamma^{ij}_{s}\Gamma^{sk}_{l}-\Gamma^{sj}_{l}\Gamma^{ik}_{s}$
and $R^{jk}_{hl}=g_{hi}R^{ijk}_{l}$. The equations (7b)–(7f) are nothing else but the Gauss–Peterson–Codazzi equations for submanifolds $M^{n}$ with a flat normal connection in a (pseudo-)Euclidean space of dimension $n+N$, where the metric $g$ plays the role of the first quadratic form of $M^{n}$, and $w_{\alpha}$ are the Weingarten operators corresponding to the field of pairwise orthogonal unit normals.
It was also proved [38] that the nonlocality type of $P$ is preserved under linear transformations of the independent variables $(t,x)$.
Finally, see [16, 31] for a discussion on how to define a Hamiltonian for the operator $P$ in the nonlocal case.
2.2 Third-order Hamiltonian operator
in the Doyle–Potëmin canonical form
The most general form of a third-order homogeneous Hamiltonian operator, in accordance with the definition given in [13], is
$\begin{split}R^{ij}=&g_{3}^{ij}\mathrm{D}_{x}^{3}+b_{3\,s}^{ij}u_{x}^{s}\mathrm{D}_{x}^{2}+[c_{3\,s}^{ij}u_{xx}^{s}+c_{3\,st}^{ij}u_{x}^{s}u_{x}^{t}]\mathrm{D}_{x}\\ &+d_{3\,s}^{ij}u_{xxx}^{s}+d_{3\,st}^{ij}u_{x}^{s}u_{xx}^{t}+d_{3\,srt}^{ij}u_{x}^{s}u_{x}^{r}u_{x}^{t}.\end{split}$ (8)
Nevertheless, in view of the computational difficulties, a minimal set of conditions on the coefficient of the above operator $R$ that are equivalent to its skew-adjointness and $[R,R]=0$ is not known.
It was independently proved in [11] and [41, 42] that there always exists a change of dependent variables that brings a Hamiltonian operator of the form (8) into the Doyle–Potëmin canonical form (4). This form has a drastically reduced set of coefficients, and the Hamiltonian property of such an operator is equivalent to
$c_{ijk}=\frac{1}{3}(f_{ik,j}-f_{ij,k}),$ (9a)
$f_{ij,k}+f_{jk,i}+f_{ki,j}=0,$ (9b)
$c_{ijk,l}=-f^{pq}c_{pil}c_{qjk}.$ (9c)
Here, $c_{ijk}=f_{iq}f_{jp}c_{k}^{pq}$ and, of course, $(f_{ij})=(f^{ij})^{-1}$.
A covariant $2$-tensor $f$ satisfying equations (9b) must be a Monge metric of a quadratic line complex (see [21] and reference therein). This is an algebraic variety in the Plücker embedding of the projective space whose affine chart is $(u^{i})$. In particular, being Monge, $f_{ij}$ is a quadratic polynomial in the field variables $u$.
It was further proved in [6, 11, 41, 42] that the tensor $f_{ij}$ can be factorised as
$f_{ij}=\phi_{\alpha\beta}\psi^{\alpha}_{i}\psi^{\beta}_{j},\quad\text{\Big{(}or, in a matrix form,}\quad f=\Psi\Phi\Psi^{\top}\Big{)}$ (10)
where $\phi$ is a constant non-degenerate symmetric matrix of dimension $n$, and
$\psi_{k}^{\gamma}=\psi_{ks}^{\gamma}u^{s}+\omega_{k}^{\gamma}$
is a non-degenerate square matrix of dimension $n$, with the constants $\psi_{ij}^{\gamma}$ and $\omega_{k}^{\gamma}$ satisfying the relations
$\psi_{ij}^{\gamma}=-\psi_{ji}^{\gamma},$
$\phi_{\beta\gamma}(\psi_{il}^{\beta}\psi_{jk}^{\gamma}+\psi_{jl}^{\beta}\psi_{ki}^{\gamma}+\psi_{kl}^{\beta}\psi_{ij}^{\gamma})=0,$
$\phi_{\beta\gamma}(\omega_{i}^{\beta}\psi_{jk}^{\gamma}+\omega_{j}^{\beta}\psi_{ki}^{\gamma}+\omega_{k}^{\beta}\psi_{ij}^{\gamma})=0.$
The interested reader can find the general expression of the Hamiltonian for the operator $R$ (provided it exists for a given system) in [23].
The system that admits the operator $R$ in a canonical form is evidently conservative. On the other hand, if a system admits such an operator, then different conservative forms thereof may not necessarily admit third-order Hamiltonian operator in a canonical form, see discussion in [17, p. 664].
2.3 Compatibility and a conjecture
The only missing part of the picture are conditions under which the operators $P$ and $R$ are compatible, i.e. their Schouten bracket vanishes: $[P,R]=0$. We use the definition of Schouten bracket in [10]:
$[A_{1},A_{2}](\psi^{1},\psi^{2},\psi^{3})=$ $\big{[}(\ell_{A_{1},\psi^{1}}(A_{2}(\psi^{2})))\psi^{3}+\text{cyclic}(\psi^{1},\psi^{2},\psi^{3})$
$+(\ell_{A_{2},\psi^{1}}(A_{1}(\psi^{2})))\psi^{3}+\text{cyclic}(\psi^{1},\psi^{2},\psi^{3})\big{]},$
where $\psi^{a}=(\psi^{a}_{i}(x,u_{\sigma}))$, $a=1,2,3$, are covectors and the square brackets on the right-hand side mean that the result is to be considered up to the image of $\mathrm{D}_{x}$ (i.e., ‘total divergencies’). In the above formula we use the linearization of an operator $A$ [7]: if, in local coordinates, one has $A(\psi)^{i}=a^{ij\sigma}\mathrm{D}_{\sigma}\psi_{j}$ then
$\ell_{A,\psi}(\phi)^{i}=\frac{\partial a^{it\sigma}}{\partial u^{s}_{\tau}}\mathrm{D}_{\sigma}\psi_{t}\mathrm{D}_{\tau}\phi^{s},\quad\phi=(\phi^{s}(x,u_{\sigma})).$
In [10], presented is an algorithm to compute a divergence-free form of the bracket $[A_{1},A_{2}]$ for a wide class of nonlocal operators (weakly nonlocal operators), to which Ferapontov-type operators belong. Based on this algorithm, in [9] software packages were developed for Mathematica, Maple and Reduce.
A minimal set of conditions that are equivalent to $[P,R]=0$ would be a big help, but so far a solution to this problem has been out of reach. An alternative is to figure out a suitable ansatz for Hamiltonian operators. Known examples of bi-Hamiltonian systems of WDVV-type [44] allow us to formulate the following conjecture.
Conjecture 4.
Let a Ferapontov-type operator $P$ parameterised by a metric $g_{ij}$ and a Doyle–Potëmin operator $R$ parameterised by a Monge metric $f_{ij}$ with the Monge decomposition $f_{ij}=\psi^{\alpha}_{i}\phi_{\alpha\beta}\psi_{j}^{\beta}$ be compatible. Then the symmetric matrix
$Q^{\alpha\beta}=\psi^{\alpha}_{i}g^{ij}\psi^{\beta}_{j}$
has entries that are second-degree polynomials in the field variables.
In turn, it means that a suitable ansatz for $g$ is
$g^{ij}=\psi_{\alpha}^{i}Q^{\alpha\beta}\psi_{\beta}^{j}\quad\text{or, in a matrix form,}\quad g=\Psi^{-1}Q(\Psi^{-1})^{\top}.$
Therefore, whenever finding a Ferapontov-type operator with a generic ansatz is not feasible, we resort to the above ansatz. Luckily, when we are able to make a general computation without resorting to an ansatz, the result is always within the ansatz, thus indicating that the ansatz might be a feature of all first-order operators $P$ which are compatible with a third-order operator $R$ as above.
Finally, note that although $(Q^{\alpha\beta})$ is quadratic in its entries, it is not a Monge metric in general.
2.4 Classification of bi-Hamiltonian equations
of WDVV-type
We classify quasilinear first-order systems of PDEs in a conservative form (1),
$u_{t}=(V^{i})_{x}=V^{i}_{s}u^{s}_{x},$
that are bi-Hamiltonian with respect to a pair of operators $P$, $R$. Here, $P$ is a Ferapontov operator of the form
$P^{ij}=g^{ij}\mathrm{D}_{x}+\Gamma^{ij}_{s}u^{s}_{x}$
$+c^{11}V^{i}_{s}u^{s}_{x}\mathrm{D}_{x}^{-1}V^{j}_{r}u^{r}_{x}+c^{12}\left(V^{i}_{s}u^{s}_{x}\mathrm{D}_{x}^{-1}u^{j}_{x}+u^{i}_{x}\mathrm{D}_{x}^{-1}V^{j}_{s}u^{s}_{x}\right)+c^{22}u^{i}_{x}\mathrm{D}_{x}^{-1}u^{j}_{x};$
Note that we used only two commuting vector fields in $P$, in view of the discussion in Subsection 2.1. The operator $R$ is a third-order homogeneous local Hamiltonian operator $R$, and can always be brought to the Doyle–Potëmin canonical form by a transformation of the dependent variables. Such a transformation neither changes the locality (resp., nonlocality) of the Ferapontov operator $P$, nor it changes the shape of the system (1).
In turn, Doyle–Potëmin operators are classified for $n=1,\dots,4$ with respect to several group actions [21, 22]:
  • the maximal group of transformations of the dependent variables that preserves the Doyle–Potëmin canonical form of $R$ — the group of affine transformations $\tilde{u}^{i}=A^{i}_{j}u^{j}+A^{i}_{0}$ with constant $A^{i}_{j}$ and $A^{i}_{0}$;
  • the group of projective reciprocal transformations that fix $t$ (as it was found in [21]),
    $\begin{split}\tilde{u}^{i}=\frac{A^{i}_{j}u^{j}+A^{i}_{0}}{\Delta},\quad\Delta=A^{0}_{j}u^{j}+A^{0}_{0},\\ \mathrm{d}\tilde{x}=\Delta\,\mathrm{d}x+(A^{0}_{i}V^{i}+A^{0}_{0})\mathrm{d}t,\qquad\mathrm{d}\tilde{t}=\mathrm{d}t;\end{split}$ (11)
  • the group of general projective reciprocal transformations [22], which is generated by a $t$-fixing projective reciprocal transformation as above and an $x\leftrightarrow t$ inversion.
It is clear that each of the above groups is a subgroup of the next one in the list. We will use the above groups to classify bi-Hamiltonian systems of WDVV-type in low dimensions.
It was proved in [23] that a first-order system of conservation laws (1) admits a third-order homogeneous Hamiltonian operator (in a canonical form (4)) if and only if
$f_{is}V^{s}_{j}=f_{js}V^{s}_{i},$ (12a)
$c_{skj}V^{s}_{i}+c_{sik}V^{s}_{j}+c_{sji}V^{s}_{k}=0,$ (12b)
$f_{ks}V^{s}_{ij}=c_{ksj}V^{s}_{i}+c_{ksi}V^{s}_{j}.$ (12c)
It is important to recall that systems of the type (1) that are Hamiltonian with respect to third-order operators are linearly degenerate (or weakly nonlinear, in another terminology) and non-diagonalisable [23].
The system (12) can be used in two ways. Given a conservative system defined by $(V^{i})$ one can determine the matrix $f$ parameterising a third-order operator $R$. Since the coefficients of $f$ are second-degree polynomials, the problem becomes algebraic in nature. On the other hand, given the matrix $f$ one can determine all the systems that are Hamiltonian with respect to the corresponding operator $R$. This problem is completely solved in [23]: we have $V^{i}=\psi^{i}_{\gamma}W^{\gamma}$, where $W^{\gamma}=\eta^{\gamma}_{s}u^{s}+\xi^{\gamma}$, and the constants $\eta^{\gamma}_{m}$ and $\xi^{\gamma}$ fulfill a certain linear algebraic system.
Therefore, given a Doyle–Potëmin operator $R$ it is an algebraic problem given by the system (12) to find all quasilinear systems in a conservative form that are Hamiltonian with respect to the operator $R$. We simplify the obtained family of systems by equivalence transformations to get a subfamily $\mathcal{S}$ of nonlinear systems, $u^{i}_{t}=V^{i}_{j}u^{j}_{x}.$ In particular, every third-order homogeneous Hamiltonian operator in dimension one can be transformed to $\mathrm{D}_{x}^{3}$ using transformations of the dependent variables only. But the quasilinear conservative systems that are Hamiltonian with respect to $\mathrm{D}_{x}^{3}$ are linear due to (12). Therefore, we omit altogether the dimension one in the classification below.
Note that also systems as above do not change their form when subject to a general projective reciprocal transformation, see [23].
Finally, we check if the family $\mathcal{S}$ admits also a Ferapontov operator $P$. It is done by direct computation, assisted by computer algebra.
We use the packages from the paper [9] in order to compute Schouten bracket $[P,R]$, with given $R$ and unknown $P$ (when possible). Namely, we collect all coefficients of the Schouten bracket as a three-vector and require them to vanish, add to this system the system of conditions under which $P$ is Hamiltonian. If we can solve this system, we have an answer.
Otherwise, we assume Conjecture 4 to hold. Finding a Ferapontov operator thus reduces to the three following steps. Firstly, we solve the commutativity condition $g^{is}V^{j}_{s}=g^{js}V^{i}_{s}$. Secondly, we solve the commutativity condition $\Gamma^{sij}V^{k}_{s}=\Gamma^{skj}V^{i}_{s}$ coming from (7b) and the fact that the system $\mathcal{S}$ is in a conservative form, where $\Gamma^{lij}=\frac{1}{2}(g^{is}g^{jl}_{,s}+g^{ls}g^{ji}_{,s}-g^{js}g^{il}_{,s})$. At this point, the matrix $Q$ and therefore $g$ is completely determined and to find the matrix $(c^{\alpha\beta})$ we solve the system (7f).
Strictly speaking, we do not have a proof of the fact that the operator $P$ preserves its shape under a general projective reciprocal transformation thus remaining of Ferapontov type. However, the transformed operator will be again homogeneous of degree $1$, and computational experiments show that it will be again of the same type. We conjecture that this is a general property.
3 Two-component systems
3.1 Affine classification
It was shown in [21] that up to affine transformations $\tilde{u}^{i}=A^{i}_{j}u^{j}+A^{i}_{0}$, $i=1$, …, $n$, there are three distinct third-order homogeneous Hamiltonian operators,
$R^{(1)}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\mathrm{D}_{x}^{3},\quad R^{(2)}=\mathrm{D}_{x}\begin{pmatrix}0&\mathrm{D}_{x}\frac{1}{u^{1}}\\ \frac{1}{u^{1}}\mathrm{D}_{x}&\frac{u^{2}}{(u^{1})^{2}}\mathrm{D}_{x}+\mathrm{D}_{x}\frac{u^{2}}{(u^{1})^{2}}\end{pmatrix}\mathrm{D}_{x},$
$R^{(3)}=\mathrm{D}_{x}\begin{pmatrix}\mathrm{D}_{x}&\mathrm{D}_{x}\frac{u^{2}}{u^{1}}\\ \frac{u^{2}}{u^{1}}\mathrm{D}_{x}&\frac{(u^{2})^{2}+1}{2(u^{1})^{2}}\mathrm{D}_{x}+\mathrm{D}_{x}\frac{(u^{2})^{2}+1}{2(u^{1})^{2}}\end{pmatrix}\mathrm{D}_{x}.$
The operator $R^{(1)}$ is admitted by linear systems of PDEs and therefore we pay no attention to it here.
The operator $R^{(2)}$ is admitted by a family of quasilinear systems
$u^{1}_{t}=(\alpha u^{1}+\beta u^{2})_{x},\quad u^{2}_{t}=\left(\alpha u^{2}+\frac{\beta(u^{2})^{2}+\gamma}{u^{1}}\right)_{x}.$
There are two inequivalent cases: $(\beta,\alpha)=(1,0)$ and $(\beta,\alpha)=(0,1)$. The second case is degenerate in a sense that the corresponding system is partially coupled and one equation can be solved explicitly, effectively making the second one linear. So, we will focus on the first case only.
Theorem 5.
The system
$u^{1}_{t}=u^{2}_{x},\quad u^{2}_{t}=\left(\frac{(u^{2})^{2}+\gamma}{u^{1}}\right)_{x}$
is Hamiltonian with respect to three first-order local homogeneous Hamiltonian operators $P^{(2,i)}$, $i=1$, $2$, $3$, parameterised by the metrics
$g^{(2,1)}=\begin{pmatrix}-u^{1}&0\\ 0&\frac{(u^{2})^{2}+\gamma}{u^{1}}\end{pmatrix},\quad g^{(2,2)}=\begin{pmatrix}0&u^{1}\\ u^{1}&2u^{2}\end{pmatrix},$
$g^{(2,3)}=\begin{pmatrix}2u^{2}&\frac{(u^{2})^{2}+\gamma}{u^{1}}\\ \frac{(u^{2})^{2}+\gamma}{u^{1}}&0\end{pmatrix}.$
The first-order operators above are mutually compatible as well as compatible with the operator $R^{(2)}$.
The operators $P^{(2,i)}$ belong to the list [28] of all first-order local homogeneous Hamiltonian operators that are compatible with $R^{(2)}$.
The operator $R^{(3)}$ is admitted by a family of quasilinear systems
$u^{1}_{t}=(\alpha u^{1}+\beta u^{2})_{x},\quad u^{2}_{t}=\left(\alpha u^{2}+\frac{\beta(u^{2})^{2}+\gamma u^{2}-\beta}{u^{1}}\right)_{x}.$
Again, there are two inequivalent cases: $(\beta,\alpha)=(1,0)$ and $(\beta,\alpha)=(0,1)$, and we will only deal with the first one.
Theorem 6.
The system
$u^{1}_{t}=u^{2}_{x},\quad u^{2}_{t}=\left(\frac{(u^{2})^{2}+\gamma u^{2}-1}{u^{1}}\right)_{x}$
is Hamiltonian with respect to three first-order local homogeneous Hamiltonian operators $P^{(3,i)}$, $i=1,2,3$, parameterised by the metrics
$g^{(3,1)}=\begin{pmatrix}-u^{1}&0\\ 0&\frac{(u^{2})^{2}+\gamma u^{2}-1}{u^{1}}\end{pmatrix},\quad g^{(3,2)}=\begin{pmatrix}0&u^{1}\\ u^{1}&2u^{2}+\gamma\end{pmatrix},$
$g^{(3,3)}=\begin{pmatrix}2u^{2}+\gamma&\frac{(u^{2})^{2}+\gamma u^{2}-1}{u^{1}}\\ \frac{(u^{2})^{2}+\gamma u^{2}-1}{u^{1}}&0\end{pmatrix}.$
The first-order operators above are mutually compatible as well as compatible with the operator $R^{(3)}$.
Again, the operators $P^{(3,i)}$ belong to the list [28] of all first-order local homogeneous Hamiltonian operators that are compatible with $R^{(3)}$. To the best of our knowledge, the multi-Hamiltonian systems in Theorems 5 and 6 are not known in the literature.
Proposition 7.
Each of the metrics $(g^{ij})_{i,j=1}^{2}$ of first-order local homogeneous Hamiltonian operators $P$ that are compatible with a third-order local homogeneous Hamiltonian operator $R$ (with Monge metric $f_{ij}=\phi_{\alpha\beta}\psi^{\alpha}_{i}\psi^{\beta}_{j}$), listed in [28], can be factorised as in Conjecture 4:
$g^{ij}=\psi^{i}_{\alpha}Q^{\alpha\beta}\psi^{j}_{\beta},$
where $Q=(Q^{\alpha\beta})$ is a symmetric matrix whose entries are quadratic polynomials of field variables.
As an example, the Monge metric $f^{(3)}$ of the operator $R^{(3)}$ admits the decomposition (10) with
$h_{3}=\begin{pmatrix}(u^{2})^{2}+1&-u^{1}u^{2}\\ -u^{1}u^{2}&(u^{1})^{2}\end{pmatrix},\quad\Psi=\begin{pmatrix}-u^{2}&1\\ u^{1}&0\end{pmatrix},\quad\Phi=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},$
while the Hamiltonian operator of Dubrovin–Novikov type parameterised by the metric $g:=c_{1}g^{(3,1)}+c_{2}g^{(3,2)}+c_{3}g^{(3,3)}$ admits the decomposition (7) with
$Q=\left(\begin{matrix}\gamma(c_{2}{-}c_{3})(u^{2})^{2}{+}(c_{1}\gamma u^{2}{-}c_{1})u^{1}{+}2c_{3}u^{2}&c_{2}(u^{1})^{2}{+}c_{1}u^{1}u^{2}{-}c_{3}(u^{2})^{2}{-}c_{3}\\ c_{2}(u^{1})^{2}{+}c_{1}u^{1}u^{2}{-}c_{3}(u^{2})^{2}{-}c_{3}&-c_{1}u^{1}+2c_{3}u^{2}+c_{3}\gamma\end{matrix}\right).$
3.2 Projective classification
It follows from the works of Agafonov and Ferapontov [2, 3, 4] that the quasilinear first-order systems that are Hamiltonian with respect to a third-order homogeneous Hamiltonian operator [23] correspond to linear line congruences. These are algebraic varieties in the Plücker embedding of the Grassmannian of lines in $\mathbb{P}^{n+1}$.
The classical works of Castelnuovo imply that all such congruences can be transformed to a single one (see [23]) when $n=2$, which implies that both operators $R^{(2)}$ and $R^{(3)}$ can be transformed to $R^{(1)}$ by a reciprocal projective transformation. Thus, there is no interesting case in dimension $2$ under the action of the above group.
More precisely, we have the following result.
Proposition 8.
Any two-component quasilinear system in a conservative form is linearisable by a reciprocal projective transformation.
It can be very difficult to find the above linearizing reciprocal projective transformation explicitly, even in this low-dimensional case. Alternatively, as a canonical case of the classification one can take a physically relevant system. Thus, the Chaplygin gas system
$u_{t}+uu_{x}+\frac{v_{x}}{v^{3}}=0,\quad v_{t}+(uv)_{x}=0,$
is known [32] to admit three first-order Dubrovin–Novikov Hamiltonian operators. Moreover, as a subcase of a polytropic gas dynamics system,
$u_{t}+uu_{x}+v^{\gamma}v_{x}=0,\quad v_{t}+(uv)_{x}=0,$
it admits [36] a nonhomogeneous third-order Hamiltonian operator. But it also admits a homogeneous one. Indeed, the diagonalised form of the Chaplygin gas system [32] is
$U_{t}=VU_{x},\quad V_{t}=UV_{x},\quad\left(\text{with}\quad U=u-\frac{1}{v},\quad V=u+\frac{1}{v}\right)$
to which the system
$u^{1}_{t}=u^{2}_{x},\quad u^{2}_{t}=\left(\frac{(u^{2})^{2}-a^{2}}{u^{1}}\right)_{x}$
is reduced with the help of the point transformation of the dependent variables
$u^{1}=\frac{2a}{U-V},\quad u^{2}=\frac{a(U+V)}{U-V},$
and the system
$u^{1}_{t}=u^{2}_{x},\quad u^{2}_{t}=\left(\frac{(u^{2})^{2}+\gamma u^{2}-1}{u^{1}}\right)_{x}$
is reduced with the help of the point transformation of the dependent variables
$u^{1}=-\frac{\sqrt{\gamma^{2}+4}}{U-V},\quad u^{2}=-\frac{\sqrt{\gamma^{2}+4}}{2}\frac{U+V}{U-V}-\frac{\gamma}{2}.$
Since both pairs of transformations preserve the homogeneity of Hamiltonian operators, the Chaplygin gas system possesses a homogeneous third-order Hamiltonian operator (which is not in Doyle–Potëmin canonical form).
4 Three-component systems
An affine classification is no longer feasible when $n>2$, due to a large amount of cases and subcases that we would have. We resort to two distinct projective classifications.
4.1 Partial projective classification
Under the action of projective reciprocal transformations that fix $t$, there are six classes of third-order homogeneous Hamiltonian operators $R^{(i)}$, see [21]. They are defined by the following Monge metrics,
$f^{(1)}=\begin{pmatrix}(u^{2})^{2}+\mu&-u^{1}u^{2}-u^{3}&2u^{2}\\ -u^{1}u^{2}-u^{3}&(u^{1})^{2}+\mu(u^{3})^{2}&-\mu u^{2}u^{3}-u^{1}\\ 2u^{2}&-\mu u^{2}u^{3}-u^{1}&\mu(u^{2})^{2}+1\end{pmatrix},$
$f^{(2)}=\begin{pmatrix}(u^{2})^{2}+1&-u^{1}u^{2}-u^{3}&2u^{2}\\ -u^{1}u^{2}-u^{3}&(u^{1})^{2}&-u^{1}\\ 2u^{2}&-u^{1}&1\end{pmatrix},\ f^{(3)}=\begin{pmatrix}(u^{2})^{2}+1&-u^{1}u^{2}&0\\ -u^{1}u^{2}&(u^{1})^{2}&0\\ 0&0&1\end{pmatrix},$
$f^{(4)}=\begin{pmatrix}-2u^{2}&u^{1}&0\\ u^{1}&0&0\\ 0&0&1\end{pmatrix},\quad f^{(5)}=\begin{pmatrix}-2u^{2}&u^{1}&1\\ u^{1}&1&0\\ 1&0&0\end{pmatrix},\quad f^{(6)}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
The corresponding quasilinear first-order systems of evolutionary PDEs have been found and described in [22, 23].
Case $R^{(6)}$.
The system which is Hamiltonian with respect to $R^{(6)}$ is linear, hence it is out of consideration here.
Case $R^{(5)}$.
The conservative quasilinear system of PDEs that is determined by the operator $R^{(5)}$ is (6), which is equivalent to a WDVV equation. The bi-Hamiltonian pair for such a system was found in [17], see Introduction.
Case $R^{(4)}$.
The system of PDEs that is determined by $R^{(4)}$ is
$u^{1}_{t}=u^{2}_{x},\quad u^{2}_{t}=\left(\frac{(u^{2})^{2}+u^{3}}{u^{1}}\right)_{x},\quad u^{3}_{t}=u^{1}_{x},$
Setting $u^{1}=f_{xxt}$, $u^{2}=f_{xtt}$, $u^{3}=f_{xxx}$ we obtain $f_{xxx}=f_{ttt}f_{xxt}-f_{xtt}^{2}$, which is equivalent to the WDVV equation (5) under the interchange of $x$ and $t$. Its bi-Hamiltonian representation by means of a compatible pair $P$ and $R^{(4)}$ as in this paper was constructed in [25, 26].
Case $R^{(3)}$.
The integrability of the system of PDEs determined by $R^{(3)}$,
$u^{1}_{t}=(u^{2}+u^{3})_{x},\quad u^{2}_{t}=\left(\frac{u^{2}(u^{2}+u^{3})-1}{u^{1}}\right)_{x},\quad u^{3}_{t}=u^{1}_{x},$
was first determined in [1]. Its third-order Hamiltonian structure was found in [23]. In [44], a criterion from [8] was used to find the metric $g$ of a first-order operator. Namely, for non-diagonalisable quasilinear systems of PDEs, the metric $g$ is proportional to a contraction of the square of the Haantjies tensor of the velocity matrix $(V^{i}_{j})$ of the system:
$g_{ij}=f(u)H^{\alpha}_{i\beta}H^{\beta}_{j\alpha}$
(see [8] for the definition of the Haantjies tensor).
What was new is that in this case the first-order homogeneous Hamiltonian operator $P$ was nonlocal, $c^{11}=c^{22}=-1$, $c^{12}=c^{21}=0$. The metric is written down in [9], and here we present it in accordance with Conjecture 4. The Monge metric $f^{(3)}$ admits a Monge decomposition $(f^{(3)}_{ij})=\Psi\Phi\Psi^{\top}$, where the metric $g=\Psi^{-1}Q(\Psi^{-1})^{\top}$, where
$Q^{11}=4(u^{1})^{2}+(u^{2})^{2}+1,\quad Q^{12}=-3u^{1},\quad Q^{13}=-2u^{2}-u^{3},$
$Q^{22}=(u^{1})^{2}+(u^{3})^{2}+4,\quad Q^{23}=u^{1}(u^{2}+2u^{3}),$
$Q^{33}=(u^{1})^{2}+(u^{2}+2u^{3})^{2}+1,$
$\Psi=(\psi^{\alpha}_{i})=\begin{pmatrix}-u^{2}&0&1\\ u^{1}&0&0\\ 0&1&0\end{pmatrix},\quad\Phi=(\phi_{\alpha\beta})=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
The system of PDEs that is determined by $R^{(3)}$ is equivalent to a particular WDVV equation [14] obtained when $\eta=\text{Id}$, the identity matrix. Setting $u^{1}=f_{xxt}$, $u^{2}=f_{xtt}-f_{xxx}$, $u^{3}=f_{xxx}$, we have
$f_{xxt}^{2}-f_{xxx}f_{xtt}+f_{xtt}^{2}-f_{xxt}f_{ttt}-1=0.$
At this point, it is natural to ask the question if in the case $n=3$ all the items in the projective classification of operators $R$ determine WDVV systems of PDEs. The answer is negative. Indeed, WDVV equations in dimension three were classified in [33], and in [44] the bi-Hamiltonian pairs for each member of the classification were found to be the pairs considered in this paper. It turned out that the projective classes of the third-order homogeneous Hamiltonian operators $R$ were that of $R^{(3)}$, $R^{(4)}$, $R^{(5)}$. It means that the systems of PDEs that are determined by $R^{(1)}$ and $R^{(2)}$ are not WDVV systems. That is one of the main motivations for this paper: there are more WDVV-type equations than WDVV systems.
Case $R^{(2)}$.
A Monge decomposition $(f^{(2)}_{ij})=\Psi\Phi\Psi^{\top}$ of the Monge metric $f^{(2)}$ is given by
$\Psi=(\psi^{\alpha}_{i})=\begin{pmatrix}u^{2}&0&1\\ -u^{1}&-u^{3}&0\\ 1&u^{2}&0\end{pmatrix},\quad\Phi=(\phi_{\alpha\beta})=\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&1\end{pmatrix}.$
The system $u^{i}_{t}=(V^{i})_{x}$ admitted by $R^{(2)}$ is (up to a transformation in the stabilizer of $R^{(2)}$ that reduces the number of constants)
$\begin{split}u^{1}_{t}=&(\alpha u^{2}+\beta u^{3})_{x},\\ u^{2}_{t}=&\left(\frac{((u^{2})^{2}-1)(\alpha u^{2}+\beta u^{3})-(\gamma+\delta u^{1})}{S}\right)_{x},\\ u^{3}_{t}=&\left(\frac{(u^{2}u^{3}-u^{1})(\alpha u^{2}+\beta u^{3})-u^{1}(\gamma+\delta u^{1})}{S}\right)_{x},\end{split}$ (13)
where $S=u^{1}u^{2}-u^{3}$ is proportional to $\sqrt{\det(f^{(2)})}$ and $\alpha,\beta,\gamma,\delta$ are arbitrary constants. The system admits a Hamiltonian formulation through $R^{(2)}$ and the nonlocal Hamiltonian
$H=\int\left(\frac{\alpha}{2}u^{3}(\mathrm{D}_{x}^{-1}u^{2})^{2}+\beta u^{3}(\mathrm{D}_{x}^{-1}u^{2})(\mathrm{D}_{x}^{-1}u^{3})-\frac{\gamma}{2}x^{2}u^{1}-\delta xu^{1}(\mathrm{D}_{x}^{-1}u^{1})\right)\mathrm{d}x.$
Again, this system is linearly degenerate, and non-diagonalisable for generic values of parameters; it is diagonalisable if and only if $\alpha\delta-\beta\gamma=0$.
Theorem 9.
The system (13) has a unique Ferapontov-type Hamiltonian operator $P$, which is compatible with $R^{(2)}$, fulfills Conjecture 4 with the metric $(g^{ij})=\Psi^{-1}Q(\Psi^{-1})^{\top}$, $c^{11}=3$, $c^{12}=c^{21}=0$, $c^{22}=-\beta^{2}$. Here,
$Q^{11}=2(A^{2}+B^{2}+4BC+2AC),\quad Q^{12}=2(3AD-BC),$
$Q^{13}=2B(2A+3C),\quad Q^{22}=-2(2A+C)(2A+3C),$
$Q^{23}=8A^{2}+10AC+2BD,\quad Q^{33}=-6A^{2}+2B^{2},$
$A=\alpha u^{2}+\beta u^{3},\quad B=\beta u^{1}+\alpha,\quad C=\delta u^{1}+\gamma,\quad D=\delta u^{3}+\gamma u^{2}.$
Case $R^{(1)}$.
In this subsection, $\mu^{2}\neq 1$. The Monge metric $f^{(1)}$ admits a Monge decomposition $(f^{(1)}_{ij})=\Psi\Phi\Psi^{\top}$, where
$(\psi^{\alpha}_{i})=\begin{pmatrix}u^{2}&0&1\\ -u^{1}&-u^{3}&0\\ 1&u^{2}&0\end{pmatrix},\quad(\phi_{\alpha\beta})=\begin{pmatrix}1&0&0\\ 0&\mu&1\\ 0&1&\mu\end{pmatrix}.$
The system $u^{i}_{t}=(V^{i})_{x}$ admitted by $R^{(1)}$ is (up to a transformation in the stabilizer of $R^{(2)}$ that reduces the number of constants) is
$\begin{split}u^{1}_{t}=&(\alpha u^{2}+\beta u^{3})_{x},\\ u^{2}_{t}=&\left(\frac{\left((u^{2})^{2}-\mu\right)(\alpha u^{2}+\beta u^{3})+\gamma(1-\mu(u^{2})^{2})+\delta(u^{1}-\mu u^{2}u^{3})}{S}\right)_{x},\\ u^{3}_{t}=&\Big{(}\frac{\alpha u^{3}\left((u^{2})^{2}-\mu\right)+\beta u^{3}(u^{2}u^{3}-\mu u^{1})}{S}\\ &\qquad+\frac{\gamma(u^{1}-\mu u^{2}u^{3})+\delta\left((u^{1})^{2}-\mu(u^{3})^{2}\right)}{S}\Big{)}_{x},\end{split}$ (14)
where $S=u^{1}u^{2}-u^{3}$ is proportional to $\sqrt{\det(f^{(1)})}$, and $\alpha,\beta,\gamma,\delta$ are arbitrary constants. Its nonlocal Hamiltonian is
$\begin{array}[]{c}H={\displaystyle\int}\Big{(}\frac{\alpha}{2}(2\mu xu^{1}\mathrm{D}_{x}^{-1}u^{2}+u^{3}(\mathrm{D}_{x}^{-1}u^{2})^{2}+\mu x^{2}u^{3})+\beta u^{3}(1-\mu^{2})\mathrm{D}_{x}^{-1}u^{2}\mathrm{D}_{x}^{-1}u^{3}\\ \\ +\delta(xu^{1}\mathrm{D}_{x}^{-1}u^{1}+\mu u^{3}\mathrm{D}_{x}^{-1}u^{1}\mathrm{D}_{x}^{-1}u^{2}+\mu u^{1}\mathrm{D}_{x}^{-1}u^{2}\mathrm{D}_{x}^{-1}u^{3}+\mu xu^{3}\mathrm{D}_{x}^{-1}u^{3})\\ \\ +\frac{1}{2}\gamma(\mu u^{1}(\mathrm{D}_{x}^{-1}u^{2})^{2}+x^{2}u^{1}+2\mu xu^{3}\mathrm{D}_{x}^{-1}u^{2}\Big{)}\mathrm{d}x.\end{array}$
The system (14) is linearly degenerate, and non-diagonalisable for generic values of parameters ($\alpha\delta-\beta\gamma\neq 0$).
Theorem 10.
The system (14) has a unique first-order nonlocal Hamiltonian operator $P$ that fulfills Conjecture 4 with the metric $(g^{ij})=\Psi^{-1}Q(\Psi^{-1})^{\top}$ and is compatible with $R^{(1)}$,
$c^{11}=\mu^{2}+3,\quad c^{12}=c^{21}=-4\mu\delta,\quad c^{22}=\mu^{3}\beta^{2}+4\mu^{2}\delta^{2}-\mu\beta^{2},$
$Q^{11}=-(\mu^{2}-1)\left(\mu^{2}(A+C)^{2}+\mu(B^{2}+D^{2})-2BD-4EF\right),$
$Q^{12}=-(\mu^{2}-1)(\mu ED-FB),\quad Q^{13}=-(\mu^{2}-1)(\mu B(2E+F)-3DE),$
$Q^{22}=-F^{2}\mu^{3}-\mu^{2}(4A^{2}+D^{2})+\mu(8BD+F^{2})-3D^{2},$
$Q^{23}=-\mu^{2}(2BD+(u^{1}u^{2}-u^{3})(\alpha\delta-\beta\gamma))+4\mu(B^{2}+D^{2})-5BD-EF,$
$Q^{33}=-\mu^{3}E^{2}-\mu^{2}(B^{2}+4D^{2})+\mu(E^{2}+8BD)-3B^{2},$
$A=\beta u^{1}+\delta u^{3},\quad B=\alpha u^{2}+\beta u^{3},\quad C=\gamma u^{2}+\alpha,$
$D=\delta u^{1}+\gamma,\quad E=\beta u^{1}+\alpha,\quad F=\delta u^{3}+\gamma u^{2}.$
4.2 General projective classification
In [23], it was proved that third-order homogeneous Hamiltonian operators in the Doyle–Potëmin canonical form are invariant also with respect to transformations that exchange $t$ and $x$. This, together with projective reciprocal transformations that fix $t$, generate a larger group of reciprocal transformations of the following type:
$\mathrm{d}\tilde{x}=(A_{i}u^{i}+A_{0})\mathrm{d}x+(A_{i}V^{i}+C_{0})\mathrm{d}t,$
$\mathrm{d}\tilde{t}=(B_{i}u^{i}+B_{0})\mathrm{d}x+(B_{i}V^{i}+D_{0})\mathrm{d}t,$
coupled with affine transformations of the dependent variables [2, 4]. Such transformations act as $\mathrm{SL}(n+2)$ transformations on the linear line congruence that corresponds to the quasilinear first-order systems of PDEs determined by third-order homogeneous Hamiltonian operators.
According to classical results by Castelnuovo, there are four distinct classes of linear line congruences in $\mathbb{P}^{4}$ under the action of $\mathrm{SL}(5)$: only two of them are endowed with third-order homogeneous Hamiltonian operator, they correspond to $R^{(5)}$ and $R^{(6)}$, see [23].
It turns out that the $6$ classes that we discussed in the previous section can be transformed to $2$ classes, one of which contains a linear system, and the other contains the simplest WDVV equation (6). It provides a proof that all systems that we discussed in the previous section are bi-Hamiltonian and integrable.
However, we stress that finding equivalence transformations between a given system or a Hamiltonian operator and a representative of an corresponding equivalence class can be extremely challenging, even with the most advanced computer algebra systems. So, having a direct proof and methods to efficiently compute systems and Hamiltonian operators proves to be an invaluable set of tools.
5 Four-component systems
There exists [22] a projective classification of third-order homogeneous Hamiltonian operators in $n=4$. The group acting on the operators is that of $t$-fixing projective reciprocal transformations.
Unfortunately, unlike cases with $n\leq 3$, not all classes of third-order homogeneous Hamiltonian operators and associated systems admit a compatible first-order local or nonlocal Hamiltonian operator that fulfill the Conjecture 4. As a distinguished example, let us consider systems of the type
$u^{1}_{t}=u^{2}_{x},\quad u^{2}_{t}=u^{3}_{x},\quad u^{3}_{t}=u^{4}_{x},\quad u^{4}_{t}=(f(u))_{x}.$ (15)
Linearly degenerate systems of the above type have been studied in [1]. In [22] it is proved that the above system is Hamiltonian with respect to a third-order homogeneous Hamiltonian operator $R$ only for two values of $f$:
$f_{1}(u)=(u^{2})^{2}-u^{1}u^{3},\qquad f_{2}(u)=(u^{3})^{2}-u^{2}u^{4}+u^{1}.$ (16)
Proposition 11.
There does not exist a matrix $(g^{ij})$ fulfilling the Conjecture 4 for the systems (15), (16).
The above proposition does not exclude the possibility that an operator $P$ that does not fulfill the Conjecture and is still compatible with the operator $R$ exists for the above system; however, given the fact that the Conjecture has been verified in a substantial number of cases (and never be disproven at the moment) we think that such a possibility has a little chance.
On the other hand, it is believed that there is a unique integrable case within the class of systems of conservation laws that admits a Hamiltonian formulation through a third-order homogeneous Hamiltonian operator (see the discussion at the end of [18]). This is represented by the system
$\begin{split}&u^{1}_{t}=u^{3}_{x},\\ &u^{2}_{t}=u^{4}_{x},\\ &u^{3}_{t}=\left(\frac{u^{1}u^{2}u^{4}+u^{3}((u^{3})^{2}+(u^{4})^{2}-(u^{2})^{2}-1)}{u^{1}u^{3}+u^{2}u^{4}}\right)_{x},\\ &u^{4}_{t}=\left(\frac{u^{1}u^{2}u^{3}+u^{4}((u^{3})^{2}+(u^{4})^{2}-(u^{1})^{2}-1)}{u^{1}u^{3}+u^{2}u^{4}}\right)_{x},\end{split}$ (17)
which is known to possess a Lax pair and a Doyle–Potëmin Hamiltonian operator $R$ parameterised by a Monge metric $f=(f_{ij})$ [23],
$(f_{ij})=\begin{pmatrix}(u^{2})^{2}{+}(u^{3})^{2}+1&-u^{1}u^{2}+u^{3}u^{4}&-u^{1}u^{3}+u^{2}u^{4}&-2u^{2}u^{3}\\ -u^{1}u^{2}+u^{3}u^{4}&(u^{1})^{2}{+}(u^{4})^{2}{+}1&-2u^{1}u^{4}&u^{1}u^{3}-u^{2}u^{4}\\ -u^{1}u^{3}+u^{2}u^{4}&-2u^{1}u^{4}&(u^{1})^{2}{+}(u^{4})^{2}&u^{1}u^{2}-u^{3}u^{4}\\ -2u^{2}u^{3}&u^{1}u^{3}-u^{2}u^{4}&u^{1}u^{2}-u^{3}u^{4}&(u^{2})^{2}{+}(u^{3})^{2}\end{pmatrix}$
The above Monge metric can be factorised as $f=\Psi\Phi\Psi^{\top}=\psi^{\alpha}_{i}\phi_{\alpha\beta}\psi^{\beta}_{j}$,
$\Psi=\begin{pmatrix}-u^{2}&-u^{3}&1&0\\ u^{1}&-u^{4}&0&1\\ -u^{4}&u^{1}&0&0\\ u^{3}&u^{2}&0&0\end{pmatrix},\quad\Phi=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}$
It is an original result the fact that the above system is bi-Hamiltonian of the type under consideration.
Theorem 12.
The system (17) is Hamiltonian with respect to a first-order nonlocal Hamiltonian operator $P$ that is compatible with $R$ and is defined by the metric $g=(g^{ij})$ splitting as in Conjecture 4, $c^{11}=c^{22}=1$, $c^{12}=c^{21}=0$. Here,
$Q^{11}=(u^{1})^{2}+(u^{2})^{2}+(u^{3})^{2}+(u^{4})^{2},\quad Q^{12}=-2u^{1}u^{4}+2u^{2}u^{3},\quad Q^{13}=-u^{2},$
$Q^{14}=u^{1},\quad Q^{22}=(u^{1})^{2}+(u^{2})^{2}+(u^{3})^{2}+(u^{4})^{2}+1,\quad Q^{23}=-2u^{3},$
$Q^{24}=-2u^{4},\quad Q^{33}=(u^{1})^{2}+(u^{3})^{2}+1,\quad Q^{34}=u^{1}u^{2}+u^{3}u^{4},$
$Q^{44}=(u^{2})^{2}+(u^{4})^{2}+1.$
6 And beyond
To the best of our knowledge, there is only one known example of a system with a WDVV-type bi-Hamiltonian structure for $n>4$. It comes from the WDVV equations in $N=4$ dimensions [39] (here, $n=6$). A new example not related with WDVV equations is given below.
In [24, Eq. (11)] the problem of finding integrable Lagrangians within a certain class is reformulated as the problem of finding solutions to two commuting quasilinear systems of first-order PDEs of the type $u^{i}_{t}=V^{i}_{j}(u^{k})u^{j}_{x}$.
Here, we rewrite the above two systems in conservative form; one of them is:
$\begin{split}&u^{1}_{t}=\left(\frac{4(u^{1})^{3}u^{3}-4(u^{1})^{2}u^{6}+2u^{1}u^{2}u^{4}-(u^{4})^{2}}{2(2u^{1}u^{2}u^{3}-u^{1}u^{5}-u^{3}u^{4})}\right)_{x},\\ &u^{2}_{t}=u^{1}_{x},\\ &u^{3}_{t}=\left(\frac{4(u^{1})^{2}(u^{3})^{2}-4u^{1}u^{3}u^{6}+u^{4}u^{5}}{2(2u^{1}u^{2}u^{3}-u^{1}u^{5}-u^{3}u^{4})}\right)_{x},\\ &u^{4}_{t}=\left(\frac{2(u^{1})^{3}u^{5}-4(u^{1})^{2}u^{2}u^{6}+2(u^{1})^{2}u^{3}u^{4}+2u^{1}(u^{2})^{2}u^{4}-u^{2}(u^{4})^{2}}{2(2u^{1}u^{2}u^{3}-u^{1}u^{5}-u^{3}u^{4})}\right)_{x},\\ &u^{5}_{t}=\left(\frac{2(u^{1})^{2}u^{3}u^{5}+2u^{1}(u^{3})^{2}u^{4}-2u^{1}u^{5}u^{6}+u^{2}u^{4}u^{5}-2u^{3}u^{4}u^{6}}{2(2u^{1}u^{2}u^{3}-u^{1}u^{5}-u^{3}u^{4})}\right)_{x},\\ &u^{6}_{t}=\frac{1}{2}u^{4}_{x}.\end{split}$ (18)
Proposition 13.
System (18) possesses a third-order Hamiltonian operator $R$ parameterised by a Monge metric $f=(f_{ij})$, where
$(f_{ij})=\begin{pmatrix}(u^{3})^{2}&-\frac{u^{5}+u^{2}u^{3}}{2}&-u^{1}u^{3}+\frac{(u^{2})^{2}}{2}&0&-\frac{u^{2}}{2}&0\\ -\frac{u^{5}+u^{2}u^{3}}{2}&u^{1}u^{3}+u^{6}&-\frac{u^{4}+u^{1}u^{2}}{2}&u^{3}&u^{1}&-\frac{u^{2}}{2}\\ -u^{1}u^{3}+\frac{(u^{2})^{2}}{2}&-\frac{u^{4}+u^{1}u^{2}}{2}&(u^{1})^{2}&-\frac{u^{2}}{2}&0&0\\ 0&u^{3}&-\frac{u^{2}}{2}&0&\frac{1}{2}&0\\ -\frac{u^{2}}{2}&u^{1}&0&\frac{1}{2}&0&0\\ 0&-\frac{u^{2}}{2}&0&0&0&1\end{pmatrix}$
Proof.
We make use of equations (12) for the above system, with an unknown Monge metric. We find the above Monge metric as the unique solution (up to constant multiples). ∎
The above Monge metric can be factorised as $f=\Psi\Phi\Psi^{\top}=\psi^{\alpha}_{i}\phi_{\alpha\beta}\psi^{\beta}_{j}$ where
$\Psi=\begin{pmatrix}-u^{3}&-u^{2}&u^{5}&0&0&0\\ 0&u^{1}&-u^{6}&u^{3}&1&0\\ u^{1}&0&u^{4}&-u^{2}&0&0\\ 0&1&-u^{3}&0&0&0\\ 0&0&-u^{1}&1&0&0\\ 0&0&u^{2}&0&0&1\end{pmatrix},\quad\Phi=\begin{pmatrix}1&0&0&0&0&0\\ 0&0&0&\frac{1}{2}&0&0\\ 0&0&0&0&-\frac{1}{2}&0\\ 0&\frac{1}{2}&0&0&0&0\\ 0&0&-\frac{1}{2}&0&0&0\\ 0&0&0&0&0&1\end{pmatrix}$
Theorem 14.
The system (18) is Hamiltonian with respect to a first-order local Hamiltonian operator $P$ that is defined by the metric $g=(g^{ij})$ splitting as in Conjecture 4 and compatible with $R$. Here,
$Q^{11}=2u^{1}u^{3},\quad Q^{12}=u^{1}u^{2}-u^{4},\quad Q^{13}=-2u^{1}u^{5}+2u^{3}u^{4},$
$Q^{14}=-u^{2}u^{3}+u^{5},\quad Q^{15}=0,\quad Q^{16}=0,\quad Q^{22}=2(u^{1})^{2},$
$Q^{23}=-4u^{1}u^{6}+2u^{2}u^{4},\quad Q^{24}=2u^{1}u^{3}-(u^{2})^{2}+4u^{6},\quad Q^{25}=4u^{1},$
$Q^{26}=2u^{4},\quad Q^{33}=-2u^{4}u^{5}+2(u^{6})^{2},\quad Q^{34}=2u^{2}u^{5}-4u^{3}u^{6},$
$Q^{35}=-4u^{1}u^{3}+(u^{2})^{2}-2u^{6},\quad Q^{36}=u^{2}u^{6}-u^{1}u^{5}-u^{3}u^{4},\quad Q^{44}=2(u^{3})^{2},$
$Q^{45}=4u^{3},\quad Q^{46}=2u^{5},\quad Q^{55}=2,\quad Q^{56}=u^{2},\quad Q^{24}=2u^{6}.$
Remark 15.
It is interesting to observe that there is another quasilinear first-order system in a conservative form that can be deduced from the two quasilinear systems in [24, Eq. (11)]. The above bi-Hamiltonian structure of WDVV-type holds also for this system, a phenomenon that has already been observed for the four-component WDVV systems in [20].
Remark 16.
A further example of bi-Hamiltonian structure of WDVV-type is incomplete, but interesting. Indeed, the simplest case of Oriented Associativity equation has a first-order local operator $P$ and a third-order nonlocal operator $R$, which suggests the possibility that the class of bi-Hamiltonian structures of WDVV type can be further enlarged.
It must be stressed that in the above case compatibility of the two operators $P$, $R$ has never been proved, also in view of the computational complexity of the problem.
Acknowledgements
We wish to express our thanks to A. Caruso, E.V. Ferapontov, P. Lorenzoni and A.C. Norman for stimulating discussions.
SO and RV acknowledge financial support from Istituto Nazionale di Fisica Nucleare (INFN) by IS-CSN4 Mathematical Methods of Nonlinear Physics (MMNLP), from the GNFM of Istituto Nazionale di Alta Matematica “F. Severi” (INdAM) and from Dipartimento di Matematica e Fisica “E. De Giorgi” of the Università del Salento. RV also acknowledges financial support from PRIN 2022 (2022TEB52W) - PE1 - The charm of integrability: from nonlinear waves to random matrices.
This work is (partially) supported by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU.